OpenStudy (devinturn):

i just cant figure it out If f(x)=1/2 x-6 what is the equation for f^–1(x)?

OpenStudy (royalranger):

f-1(x)? do you mean f(-x)?

OpenStudy (devinturn):

no i mean inverse!!!!!!!!!! :(((

OpenStudy (devinturn):

please plese plese plez :((( sad boy

OpenStudy (mathmale):

I'd suggest you write \[f ^{-1}(x)\]

OpenStudy (mathmale):

for "inverse function of x." Otherwise your presentation is subject to being misinterpreted.

OpenStudy (mathmale):

also, your "1/2" in f(x)=1/2 x-6 could be misinterpreted. Write it as (1/2) or as

OpenStudy (mathmale):

If f(x)=(1/2)x - 6, please replace f(x) with y. What do you get?

OpenStudy (mathmale):

Welcome to OpenStudy!

OpenStudy (mathmale):

Devin, if you want ongoing help, you may have to stick around longer after having posted a question. Otherwise you may lose the people who have started helping you or were considering helping you.